OpenStudy (anonymous):

y=(-x)/(x^2+1) as x->infinity, y=???? why

OpenStudy (anonymous):

Since \(x^2\) dominates \(x\) the limit as \(x\) approaches infinity is 0

OpenStudy (anonymous):

\[-\frac{x}{x^2+1}= -\frac{1}{\displaystyle\frac{x^2}{x}+\frac{1}{x}} = -\frac{1}{x+\displaystyle\frac{1}{x}}\]\[\lim_{x \to \infty}-\frac{1}{x+\displaystyle\frac{1}{x}} = 0\]

OpenStudy (anonymous):

why divide x by x square

OpenStudy (anonymous):

It is to demonstrate that the denominator dominates the numerator as x approaches infinity.

OpenStudy (anonymous):

huh...?? is it we must always make the top a constant?

OpenStudy (anonymous):

What happens to \(\displaystyle\frac{1}{x}\) for big \(x\)

OpenStudy (anonymous):

1/0

OpenStudy (anonymous):

i mean zero

OpenStudy (anonymous):

As x gets very large 1/x gets very small. As you can see above, 1/x on the bottom will disappear, but you will still have -1/x which will approach 0 as x gets large.

OpenStudy (anonymous):

so u look at things part by part? cant look at it straight?